# Stochastic theta methods for random periodic solution of stochastic differential equations under non-globally Lipschitz conditions11footnotemark: 1

### v1

@@ 2 Assumptions and random periodic solutions of SDEs
Now we can prove the existence and uniqueness of random periodic solution for (1.1), whose proof is analogous to that of [4, Theorem 2.4] and hence is omitted.
@@
Suppose that Assumption 2.1 holds. Then for any initial value $\xi$ satisfying (A4), the considered equation (1.1) admits a unique random periodic solution $X_{t}^{*}(\cdot)\in L^{2}(\Omega),t\geq 0$, given by
@@ 3 Random periodic solutions of ST methods
$\langle x-y,f(t,x)-f(t,y)\rangle\leq L_{f}|x-y|^{2},\quad x,y\in\mathbb{R}^{d},t\in\mathbb{R},$
@@
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
@@
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}+2\Delta\widetilde{C}$
@@
$~{}+(\theta^{2}-2\theta+1)\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}+2\Delta\widetilde{C}.$
@@
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}\Big{)}.$
@@
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-A\xi+f(0,\xi)|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}\bigg{)}$
@@
$\leq$ $~{}\max\bigg{\{}\frac{1+2(\lambda-\widetilde{L_{f}})\theta}{2(\lambda-\widetilde{L_{f}})\theta},~{}\frac{(p^{*}-1)\theta}{p^{*}-2},~{}\frac{\theta^{2}}{2\theta-1}\bigg{\}}:=\bar{C},$
@@
Let $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}$ and $\{\hat{Y}_{-k\tau+t_{j}}^{-k\tau}\}$ be two solutions of the BEM scheme (1) with different initial values $\xi$ and $\eta$, respectively. Suppose that Assumption 2.1 holds for both initial values $\xi$ and $\eta$. Then there exist constants $C>0$ and $C_{\Delta}\in[0,1)$, independent of $j$ and $k$, such that
$\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}-\hat{Y}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}\leq CC_{\Delta}^{j}\big{(}\mathbb{E}[|\xi-\eta|^{2}]+\mathbb{E}[|\xi-\eta|^{\gamma}]\big{)},$ (3.3)
@@
$=$ $~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|(-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau}))|^{2}\big{]}$
@@
Suppose that Assumption 2.1 holds. Then for any initial value $\xi$ satisfying (A4), the considered method (1) admits a unique random period solution $\hat{X}_{t}^{*}(\cdot)\in L^{2}(\Omega;\mathbb{R}^{d}),t\geq 0$ such that
@@ 4 Mean square convergence order of ST methods
In this section, we will establish the mean square convergence order of ST methods. To this end, the Hölder continuity of the exact solution is needed.
Lemma 4.1.
Suppose that Assumption 2.1 holds. Then for any $q\in[2,\frac{p^{*}}{\gamma}]$, $k\in\mathbb{N}$ and $s,t\geq-k\tau$, there exists a constant $C>0$, independent of $k,s,t$, such that
$\mathbb{E}\big{[}|X_{t}^{-k\tau}-X_{s}^{-k\tau}|^{q}\big{]}\leq C\big{(}|t-s|^{\frac{q}{2}}+|t-s|^{q}\big{)}.$ (4.1)
Proof.
Without loss of generality, we assume $s\leq t$ and get
$~{}\mathbb{E}\big{[}|X_{t}^{-k\tau}-X_{s}^{-k\tau}|^{q}\big{]}$
$=$ $~{}\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau}){\,\rm{d}}{r}+\int_{s}^{t}g(r,X_{r}^{-k\tau}){\,\rm{d}}{W_{r}}\Big{|}^{q}\bigg{]}$ (4.2)
$\leq$ $~{}C\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau}){\,\rm{d}}{r}\Big{|}^{q}\bigg{]}+\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}g(r,X_{r}^{-k\tau}){\,\rm{d}}{W_{r}}\Big{|}^{q}\bigg{]}.$
By the Hölder inequlity, (2.8) and Lemma 2.2, we have
$~{}\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau}){\,\rm{d}}{r}\Big{|}^{q}\bigg{]}$
$\leq$ $~{}(t-s)^{q-1}\int_{s}^{t}\mathbb{E}\big{[}|-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau})|^{q}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{q-1}\int_{s}^{t}\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{q}\big{]}+\mathbb{E}\big{[}|f(r,X_{r}^{-k\tau})|^{q}\big{]}{\,\rm{d}}{r}$ (4.3)
$\leq$ $~{}C(t-s)^{q-1}\int_{s}^{t}1+\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{q}\big{]}+\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{q\gamma}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{q}.$
Applying [14, Theorem 7.1 in Chapter 1], (2.8) and Lemma 2.2 yields
$\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}g(r,X_{r}^{-k\tau}){\,\rm{d}}{W_{r}}\Big{|}^{q}\bigg{]}\leq$ $~{}C(t-s)^{\frac{q}{2}-1}\int_{s}^{t}\mathbb{E}\big{[}|g(r,X_{r}^{-k\tau})|^{q}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{\frac{q}{2}-1}\int_{s}^{t}1+\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{\frac{q(\gamma+1)}{2}}\big{]}{\,\rm{d}}{r}$ (4.4)
$\leq$ $~{}C(t-s)^{\frac{q}{2}}.$
Substituting (4) and (4) into (4) enables us to derive (4.1). ∎
Next, we define
@@
Now for any $s\in[-k\tau+t_{j},-k\tau+t_{j+1}]$, using (2.3), (2.7) and Lemma 4.1 indicates
$~{}\mathbb{E}\big{[}|(f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|(f(s,X_{s}^{-k\tau})-f(-k\tau+t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
@@
Repeating the techniques used in (4) and exploiting (4), (4.9) as well as (4.10) implies the second part of (4.6). Thus we complete the proof. ∎
@@
where $\gamma\geq 1$ comes from (2.3).
@@
Suppose that Assumptions 2.1 and 4.3 hold. Then for any $k,j\in\mathbb{N}$, there exists a constant $C>0$, independent of $k,j$, such that
@@
$~{}+\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+2\mathbb{E}\big{[}\big{\langle}\big{(}g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}\big{)}\Delta W_{-k\tau+t_{j}}^{-k\tau},\mathcal{R}_{k,j}\big{\rangle}\big{]}$ (4.16)
@@
$=$ $~{}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
@@
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
@@
$~{}2\mathbb{E}\big{[}\big{\langle}\big{(}g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}\big{)}\Delta W_{-k\tau+t_{j}}^{-k\tau},\mathcal{R}_{j}\big{\rangle}\big{]}$
@@
$~{}+(1+\kappa\Delta)(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
@@
$\overline{C_{\Delta}}:=\max\bigg{\{}1-\frac{(\lambda-L_{f})\Delta}{1+2(\lambda-L_{f})\theta\Delta},~{}1-\frac{1}{2\theta},~{}1-\frac{2\theta-1}{2\theta^{2}}\bigg{\}}\in(0,1)$ (4.23)
@@
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}-\frac{C\Delta^{2}}{1-\overline{C_{\Delta}}}\bigg{)}$

### v2

@@ 2 Assumptions and random periodic solutions of SDEs
The forthcoming result presents the Hölder continuity of the exact solution with respect to the final time.
Lemma 2.4.
Suppose that Assumption 2.1 holds. Then for any $p\in[2,\frac{p^{*}}{\gamma}]$, $t_{0}\in\mathbb{R}$ and $s,t\geq t_{0}$, there exists a constant $C>0$, independent of $t_{0},s,t$, such that
$\mathbb{E}\big{[}|X_{t}^{t_{0}}-X_{s}^{t_{0}}|^{p}\big{]}\leq C\big{(}|t-s|^{\frac{p}{2}}+|t-s|^{p}\big{)}.$ (2.14)
Proof.
Without loss of generality, we assume $s\leq t$ and get
$~{}\mathbb{E}\big{[}|X_{t}^{t_{0}}-X_{s}^{t_{0}}|^{p}\big{]}$
$=$ $~{}\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{t_{0}}+f(r,X_{r}^{t_{0}}){\,\rm{d}}{r}+\int_{s}^{t}g(r,X_{r}^{t_{0}}){\,\rm{d}}{W_{r}}\Big{|}^{p}\bigg{]}$ (2.15)
$\leq$ $~{}C\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{t_{0}}+f(r,X_{r}^{t_{0}}){\,\rm{d}}{r}\Big{|}^{p}\bigg{]}+C\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}g(r,X_{r}^{t_{0}}){\,\rm{d}}{W_{r}}\Big{|}^{p}\bigg{]}.$
By the Hölder inequality, (2.8) and Lemma 2.2, we have
$~{}\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{t_{0}}+f(r,X_{r}^{t_{0}}){\,\rm{d}}{r}\Big{|}^{p}\bigg{]}$
$\leq$ $~{}(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\big{[}|-AX_{r}^{t_{0}}+f(r,X_{r}^{t_{0}})|^{p}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\big{[}|X_{r}^{t_{0}}|^{p}\big{]}+\mathbb{E}\big{[}|f(r,X_{r}^{t_{0}})|^{p}\big{]}{\,\rm{d}}{r}$ (2.16)
$\leq$ $~{}C(t-s)^{p-1}\int_{s}^{t}1+\mathbb{E}\big{[}|X_{r}^{t_{0}}|^{p}\big{]}+\mathbb{E}\big{[}|X_{r}^{t_{0}}|^{p\gamma}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{p}.$
Applying [14, Theorem 7.1 in Chapter 1], (2.8) and Lemma 2.2 yields
$\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}g(r,X_{r}^{t_{0}}){\,\rm{d}}{W_{r}}\Big{|}^{p}\bigg{]}\leq$ $~{}C(t-s)^{\frac{p}{2}-1}\int_{s}^{t}\mathbb{E}\big{[}|g(r,X_{r}^{t_{0}})|^{p}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{\frac{p}{2}-1}\int_{s}^{t}1+\mathbb{E}\big{[}|X_{r}^{t_{0}}|^{\frac{p(\gamma+1)}{2}}\big{]}{\,\rm{d}}{r}$ (2.17)
$\leq$ $~{}C(t-s)^{\frac{p}{2}}.$
Substituting (2) and (2) into (2) enables us to derive (2.14). ∎
Furthermore, one can show the Hölder continuity of the exact solution with respect to the initial time.
Lemma 2.5.
Suppose that Assumption 2.1 holds. Then for any $p\in[2,\frac{p^{*}}{\gamma}]$, $t\geq t_{0}$ and $t\geq s_{0}$, there exists a constant $C>0$, independent of $t,s_{0},t_{0}$, such that
$\mathbb{E}\big{[}|X_{t}^{s_{0}}(\xi)-X_{t}^{t_{0}}(\xi)|^{p}\big{]}\leq C(|t_{0}-s_{0}|^{\frac{p}{2}}+|t_{0}-s_{0}|^{p}).$
Proof.
Without loss of generality, we assume $s_{0}\leq t_{0}$ and use (2.11) to get
$X_{t}^{s_{0}}(\xi)-X_{t}^{t_{0}}(\xi)=u_{s_{0},t}(\xi)-u_{t_{0},t}(\xi)=u_{t_{0},t}(u_{s_{0},t_{0}}(\xi))-u_{t_{0},t}(\xi)=X_{t}^{t_{0}}(X_{t_{0}}^{s_{0}}(\xi))-X_{t}^{t_{0}}(\xi)$ (2.18)
almost surely. It follows from Lemmas 2.2 and 2.3, $L_{f}\in(0,\lambda)$ and $t_{0}\leq t$ that
$\mathbb{E}\big{[}|X_{t}^{s_{0}}(\xi)-X_{t}^{t_{0}}(\xi)|^{p}\big{]}=$ $~{}\mathbb{E}\big{[}|X_{t}^{t_{0}}(X_{t_{0}}^{s_{0}}(\xi))-X_{t}^{t_{0}}(\xi)|^{p}\big{]}$
$\leq$ $~{}e^{p(L_{f}-\lambda)(t-t_{0})}\mathbb{E}\big{[}|X_{t_{0}}^{s_{0}}(\xi)-\xi|^{p}\big{]}$
$\leq$ $~{}\mathbb{E}\big{[}|X_{t_{0}}^{s_{0}}(\xi)-X_{s_{0}}^{s_{0}}(\xi)|^{p}\big{]},$
which together with Lemma 2.4 finishes the proof. ∎
Now we can prove the existence and uniqueness of random periodic solution for (1.1).
@@
Suppose that Assumption 2.1 holds. Then for any initial value $\xi$ satisfying (A4), the considered equation (1.1) admits a unique random periodic solution $X_{t}^{*}\in L^{2}(\Omega;\mathbb{R}^{d}),t\geq 0$, given by
@@
Proof.
Noting that (1.1) admits a unique global solution, we will first follow the procedures presented in the proof of [24, Theorem2.5 and Proposition 5.2] to attest that (1.1) admits a global semiflow.Actually, applying Lemmas 2.3, 2.4 and 2.5 shows that for any $T_{1},T_{2}\in\mathbb{R}$ with $T_{1}<T_{2}$ and $p\in[2,\frac{p^{*}}{\gamma}]$, there exists $C(T_{1},T_{2})>0$ such that for any $x,y\in\mathbb{R}^{d}$ and $s_{0},s,t_{0},t\in[T_{1},T_{2}]$ with $s_{0}\leq s,t_{0}\leq t$,
$\mathbb{E}\big{[}|X_{s}^{s_{0}}(x)-X_{t}^{t_{0}}(y)|^{p}\big{]}\leq C(T_{1},T_{2})(|s_{0}-t_{0}|^{\frac{p}{2}}+|s-t|^{\frac{p}{2}}+|x-y|^{p}).$ (2.20)
Then $\frac{4+d}{p}<1$ and the Kolmogorov continuity theorem [25, Theorem 1.4.1] ensure that there exists a modification $v_{s,t}$ of $u_{s,t}$ in (2.10) such that $(s,t,x)\mapsto v_{s,t}(x)$ is continuous for all $\omega\in\Omega$.
Let $\{r_{n}\}_{n\in\mathbb{N}}$ be the sequence of rational numbers in $\mathbb{R}$. Then for each $n\in\mathbb{N}$, the fact $v_{s,s}(x)=u_{s,s}(x)=x$ almost surely for all $s\in\mathbb{R}$ and $x\in\mathbb{R}^{d}$ shows that there exists $\Omega_{n}\subset\Omega$ with $\mathbb{P}(\Omega_{n})=0$ such that $v_{r_{n},r_{n}}(\omega,x)=x$ for all $x\in\mathbb{R}^{d}$ and $\omega\in\Omega_{n}^{c}$. Putting $\widetilde{\Omega}:=\cup_{n=1}^{\infty}\Omega_{n}$ yields $\mathbb{P}(\widetilde{\Omega})=0$ and $v_{r_{n},r_{n}}(\omega,x)=x,n\in\mathbb{N}$ for all $x\in\mathbb{R}^{d}$ and $\omega\in\widetilde{\Omega}^{c}$. Together with the continuity of $(s,x)\mapsto v_{s,s}(\omega,x)$ for all $\omega\in\Omega$, we see that $v_{s,s}(\omega,x)=x$ for all $s\in\mathbb{R}$, $x\in\mathbb{R}^{d}$ and $\omega\in\widetilde{\Omega}^{c}$. Redefining $v_{s,t}(\omega,x):=x$ for any $s,t\in\mathbb{R}$ with $s\leq t$, $x\in\mathbb{R}^{d}$ and $\omega\in\widetilde{\Omega}$ enables us to derive $v_{s,s}(\omega,x)=x$ for all $s\in\mathbb{R}$, $x\in\mathbb{R}^{d}$ and $\omega\in\Omega$.
Observing (2.11), the uniqueness of global solution of (1.1) ensures that
$v_{s,t}(x)=v_{r,t}(v_{s,r}(x)),\quad s\leq r\leq t,~{}x\in\mathbb{R}^{d},~{}\text{almost surely}.$ (2.21)
Because the functions on the both sides of (2.21) are continuous with respect to $(s,r,t,x)$, there exists $\widehat{\Omega}\subset\Omega$ with $\mathbb{P}(\widehat{\Omega})=0$ such that $v_{s,t}(\omega,x)=v_{r,t}(\omega,v_{s,r}(\omega,x))$ for all $s,r,t\in\mathbb{R}$ with $s\leq r\leq t$, all $x\in\mathbb{R}^{d}$ and all $\omega\in\widehat{\Omega}^{c}$. Redefining $v_{s,t}(\omega,x):=x$ for any $s,t\in\mathbb{R}$ with $s\leq t$, $x\in\mathbb{R}^{d}$ and $\omega\in\widehat{\Omega}$ results in $v_{s,t}(\omega,x)=v_{r,t}(\omega,v_{s,r}(\omega,x))$ for all $s,r,t\in\mathbb{R}$ with $s\leq r\leq t$, all $x\in\mathbb{R}^{d}$ and all $\omega\in\Omega$.
Applying (A5) in Assumption 2.1 and [24, Definition 1.1] indicates that $v$ is a global semiflow for (1.1). The remainder of this proof is analogous to that of [4, Theorem 2.4] and thus omitted. ∎
@@ 3 Random periodic solutions of ST methods
$\langle x-y,f(t,x)-f(t,y)\rangle\leq L_{f}|x-y|^{2},\quad x,y\in\mathbb{R}^{d},t\in\mathbb{R}.$
@@
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j}}^{-k\tau}-f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
@@
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j}}^{-k\tau}-f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}+2\Delta\widetilde{C}$
@@
$~{}+(\theta^{2}-2\theta+1)\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j}}^{-k\tau}-f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}+2\Delta\widetilde{C}.$
@@
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j}}^{-k\tau}-f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}\Big{)}.$
@@
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|A\xi-f(0,\xi)|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}\bigg{)}$
@@
$\leq$ $~{}\max\bigg{\{}\frac{1+2(\lambda-\widetilde{L_{f}})\theta}{2(\lambda-\widetilde{L_{f}})\theta},~{}{\frac{2(p^{*}-1)\theta}{p^{*}-2}},~{}\frac{\theta^{2}}{2\theta-1}\bigg{\}}:=\bar{C},$
@@
Let $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}$ and $\{\hat{Y}_{-k\tau+t_{j}}^{-k\tau}\}$ be two solutions of the ST methods (1) with different initial values $\xi$ and $\eta$, respectively. Suppose that Assumption 2.1 holds for both initial values $\xi$ and $\eta$. Then there exists a constant $C>0$, independent of $j$ and $k$, such that
$\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}-\hat{Y}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}\leq CC_{\Delta}^{j}\Big{(}\mathbb{E}[|\xi-\eta|^{2}]+\big{(}\mathbb{E}\big{[}|\xi-\eta|^{2\gamma}\big{]}\big{)}^{\frac{1}{\gamma}}\Big{)},$ (3.3)
@@
$=$ $~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
@@
Suppose that Assumption 2.1 holds. Then for any initial value $\xi$ satisfying (A4), the considered method (1) admits a unique random period solution $\hat{X}_{t}^{*}\in L^{2}(\Omega;\mathbb{R}^{d}),t\geq 0$ such that
@@ 4 Mean square convergence order of ST methods
In this section, we will establish the mean square convergence order of ST methods. To this end, let us define
@@
Now for any $s\in[-k\tau+t_{j},-k\tau+t_{j+1}]$, using (2.3), (2.7), Lemmas 2.2 and 2.4 indicates
$~{}\mathbb{E}\big{[}|f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|f(s,X_{s}^{-k\tau})-f(-k\tau+t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
@@
Repeating the techniques used in (4) and exploiting (4), (4.5) as well as (4.6) imply the second part of (4.2). Thus we complete the proof. ∎
@@
where $\gamma\geq 1$ comes from (2.3), and $\frac{\partial f}{\partial x}$ denotes the partial derivative of $f$ with respect to the state variable $x$.
@@
Suppose that Assumptions 2.1 and 4.2 hold. Then for any $k,j\in\mathbb{N}$, there exists a constant $C>0$, independent of $k,j$, such that
@@
$~{}+\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+2\mathbb{E}\big{[}\big{\langle}\big{(}g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}\big{)}{\Delta W_{-k\tau+t_{j}}},\mathcal{R}_{k,j}\big{\rangle}\big{]}$ (4.12)
@@
$=$ $~{}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
@@
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
@@
$~{}2\mathbb{E}\big{[}\big{\langle}\big{(}g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}\big{)}{\Delta W_{-k\tau+t_{j}}},\mathcal{R}_{k,j}\big{\rangle}\big{]}$
@@
$~{}+(1+\kappa\Delta)(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
@@
$\overline{C_{\Delta}}:=\max\bigg{\{}1-\frac{(\lambda-L_{f})\Delta}{1+2(\lambda-L_{f})\theta\Delta},~{}1-\frac{1}{2\theta},~{}1-\frac{2\theta-1}{2\theta^{2}}\bigg{\}}\in{[0,1)}$ (4.19)
@@
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}-\frac{C\Delta^{2}}{1-\overline{C_{\Delta}}}\bigg{)}.$
@@
At this stage, it is necessary to compare our convergence result in Lemma 4.4 with that in [26], where the authors establish the mean square convergence rates of ST methods with $\theta\in[1/2,1]$ on finite time interval for SDEs under a coupled monotonicity condition. In contrast, the convergence result for ST methods with critical parameter $\theta=\frac{1}{2}$ does not include in Lemma 4.4, which can be attributed to the infiniteness of time interval and the superlinear growth of coefficients (see [21] for SDEs with global Lipschitz coefficients). Actually, the upper bound on the convergence error of ST method with $\theta=\frac{1}{2}$ grows exponentially with the length of time interval, making a convergence analysis on infinite time interval impossible.